# Audio Array-Based 3D UAV Trajectory Estimation with LiDAR Pseudo-Labeling

### v2

@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAV) [1] have become widely accessible, raising significant security concerns such as privacy violations [2], unauthorized surveillance [3], and risks in restricted airspace [4], including threats to civilian aviation. Their use in cross-border drug smuggling [5] and military operations underscores the urgent need for effective UAV tracking technologies [6].
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [7] [8] [9] [10] for 2D tracking or 3D detection from RADAR/LiDAR [11] [12, 13] [14]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
     The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [15, 16] in UAV trajectory estimation[17, 18]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras [19], there is a significant opportunity to develop innovative 3D tracking solutions that leverage audio information.
@@ 2.3 Gaussian Process Smoothing
$\mathbf{f}_{*}\mid\mathbf{t},\mathbf{y},t_{*}\sim\mathcal{N}(\mathbf{\mu}_{*},\mathbf{\Sigma}_{*})$
@@
$\mathbf{\mu}_{*}=k(t_{*},\mathbf{t})[k(\mathbf{t},\mathbf{t})+\sigma_{n}^{2}I]^{-1}\mathbf{y}$
@@ 2.4 Trajectory Estimation Head
$\text{Loss}=\frac{(1-\alpha)}{N}\sum_{i=1}^{N}\left({\text{P}}^{(i)}-{\xi}^{(i)}\right)^{2}+\frac{\alpha}{N}\sum_{i=1}^{N}\left({\text{P}}^{(i)}-{\Xi}^{(i)}\right)^{2}$ (1)

### v3

@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAV) [1] have become widely accessible [2], raising significant security concerns such as privacy violations [3, 4], unauthorized surveillance [5, 6], and risks in restricted airspace [7, 8, 9], including threats to civilian aviation. Their use in cross-border drug smuggling [10] and military operations [11] underscores the urgent need for effective UAV tracking technologies [12].
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [13, 14, 15, 16, 17] for 2D tracking or 3D detection from RADAR/LiDAR [18, 19] [20, 21, 4] [22]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
     The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [23, 24, 25] in UAV trajectory estimation[26, 27]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras [28], there is a significant opportunity to develop innovative 3D tracking solutions [29, 30, 31, 32, 33] that leverage audio information.
@@ 2.3 Gaussian Process Smoothing
$\mathbf{f}_{*}\mid\mathbf{t},\mathbf{y},t_{*}\sim\mathcal{N}(\mathbf{\mu}_{*},\mathbf{\Sigma}_{*}),$
@@
$\mathbf{\mu}_{*}=k(t_{*},\mathbf{t})[k(\mathbf{t},\mathbf{t})+\sigma_{n}^{2}I]^{-1}\mathbf{y},$
@@ 2.4 Trajectory Estimation Head
$\text{Loss}=\frac{(1-\alpha)}{N}\sum_{i=1}^{N}\left({\text{P}}^{(i)}-{\xi}^{(i)}\right)^{2}+\frac{\alpha}{N}\sum_{i=1}^{N}\left({\text{P}}^{(i)}-{\Xi}^{(i)}\right)^{2}$